A delivery truck is transporting boxes of two sizes: Large and small. The combined weight of a large box and a small box is 65 pounds. The truck is transporting 60 large boxes and 50 small boxes. If the truck is carrying a total of 3700 pounds in boxes, how much does each type of box weigh?
Large box: ____ pounds
Small box: _____ pounds

Answer by mananth(16949)

large box x lbs
small box y lbs
Equation for tickets total distance

1 x + 1 y = 65 .............1

Equation for amount spent
60 x + 50 y = 3700 .............2
Eliminate y
multiply (1)by -50.00
Multiply (2) by 1
-50.00 x -50.00 y = -3250.00
60.00 50.00 y = 3700.00
Add the two equations
10.00 x = 450.00
/ 10.00
x = 45.00
plug value of x in (1)
1.00 x + 1.00 y = 65.00
45.00 + 1.00 y = 65.00
1.00 y = 20.00
y= 20.00

large box 45.00 lbs
small box 20.00 lbs
m.ananth@hotmail.ca